# LEASE: Offline Preference-based Reinforcement Learning with High Sample Efficiency

### v2

@@ I Introduction
Reinforcement learning (RL) has been widely applied in robot control [1, 10] and healthcare [12, 11] fields. However, in the real-world scenarios, RL faces two serious challenges: 1) Since RL learns the optimal policy through trial and error, it is dangerous for agent to interact directly with environment, especially human-in-loop control [39]; 2) Due to the complexity of human cognitive and control system, it is difficult to design a quantitative reward function to accurately reflect the human intention or preference [9]. Much research has been conducted to solve the above two challenges.
@@ II Related Works
This section presents a brief overview of the algorithms and related theories of offline PbRL, and discusses the differences and advantages between our work and the previous related works.
The algorithm for offline PbRL. Offline PbRL eliminates the demand for interaction with the environment and handcrafted-designed reward. OPAL [34] is the first algorithm that combines offline RL and PbRL. PT [17] utilized a transformer-based architecture to design a preference model capable of generating non-Markovian rewards. OPPO [30] directly optimized the policy in a high-level embedding space without learning a separate reward function. However, the above methods demand a large amount of preference data. FTB [33] trains a diffusion model to achieve data augmentation, but this consumes a larger training cost. Instead, our method LEASE trains a simple transition model to generate unlabeled data and designs a selection mechanism to ensure the quality of the generated data, which can achieve superior performance under fewer datasets and less time.
@@ III-A Preliminaries for RL
Preference-based reinforcement learning. Different from the standard RL setting, the reward function is not available in PbRL. Instead, PbRL learns the reward function $\widehat{R}$ from preferences between pairs of trajectory segments to align the human intention [15], where a trajectory segment $\sigma$ of length $L$ is defined as a sequence of states and actions $\{s_{k},a_{k},...,s_{k+L-1},a_{k+L-1}\}\in(\mathcal{S}\times\mathcal{A})^{L}$. Given a pair of segments $(\sigma_{0},\sigma_{1})$, human choose which segment is preferred, i.e., $y\in\{0,1\}$ [14]. The preference label $y=1$ indicates $\sigma_{1}\succ\sigma_{0}$ and $y=0$ indicates $\sigma_{0}\succ\sigma_{1}$, where $\sigma_{i}\succ\sigma_{j}$ denotes that the segment $i$ is preferable to the segment $j$.
@@
After obtaining the reward function $\widehat{R}$, offline PbRL optimizes the policy by maximizing the expected discounted return $J(\pi,R)$ like standard offline RL.
@@
In practical implantation, the transition model is approximated by Gaussian distribution and the MLE method is employed to train $N_{T}$ ensemble transition models $\{T^{i}_{\varphi}=\mathcal{N}(\mu^{i}_{\varphi},\sigma^{i}_{\varphi})\}_{i=1}^{N_{T}}$ [37]. The samples are generated through $H$-step rollouts. Here, we use the trained transition to generate more trajectories to achieve data augmentation.
@@ III-B Problem Formulation
This assumption states that there must exist a reward function $R\in\mathcal{R}$ which can well-approximate the optimal reward function when the preference data can cover all data space. However, preference data tend to be limited, even fewer than a hundred in real-world scenarios. Therefore, different from previous PbRL [3, 35, 33], this paper aims to study how to learn the better reward function from the limited dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$. To broaden the coverage of preference data, motivated by model-based offline RL methods, we train the transition model through Eq. (4) to generate more unlabeled preference data $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$. Then, the pseudo labels $\widehat{y}$ for unlabeled dataset $\mathcal{D}_{u}$ can be obtained through reward model trained on $\mathcal{D}_{l}$ [3], that is
$\widehat{y}\left(\sigma_{0}^{u},\sigma_{1}^{u}\right)=\mathbbm{1}\big[P\left(\sigma_{1}^{u}\succ\sigma_{0}^{u};{\psi}\right)>0.5\big],$ (5)
where $\mathbbm{1}(\cdot)$ is indicator function. The reward function can be updated through the collected labeled dataset $\mathcal{D}_{l}$ and the generated unlabeled dataset $\mathcal{D}_{u}$. Then, the reward model can be optimized by minimizing
@@ IV Reward Learning
The reward learning involves two stages: pretraining based on collected labeled data $\mathcal{D}_{l}$ and updating based on $\mathcal{D}_{l}$ and unlabeled data $\mathcal{D}_{u}$ during policy learning, which is essentially semi-supervised learning [26]. This section focuses on designing a data selection function $f(\sigma_{0}^{u},\sigma_{1}^{u})$ to ensure the quality of generated preference data, and explaining the factors that influence the generalization ability of the reward model.
@@ IV-A Uncertainty-aware Pseudo-label Selection
where $\kappa_{p}$ and $\kappa_{\tau}$ are thresholds of confidence and uncertainty respectively. $f(\sigma_{0}^{u},\sigma_{1}^{u})=1$ denotes the generated data $(\sigma_{0}^{u},\sigma_{1}^{u})$ is selected, and $f(\sigma_{0}^{u},\sigma_{1}^{u})=0$ denotes the data is not selected.
Intuitively, more rigorous data selection criteria, characterized by higher confidence and lower variance, lead to a reduction in the number of qualified samples. While such a strategy improves the quality of the generated data, it could lead to a narrower data distribution, which might adversely affect the precision of the trained reward model. Therefore, appropriate selection of parameters $\kappa_{p}$ and $\kappa_{\tau}$ is required to enhance the performance of the reward model. Then, combining Eq. (6) and the designed selection mechanism, the reward model can be optimized by minimizing
@@
Note that the bound of pseudo-label error rate $\eta$ of the generated data filtered by the proposed selection mechanism is significantly lower than that of the unfiltered data, as the mechanism effectively eliminates low-confidence and high-uncertainty samples. This observation is experimentally validated in Table II and Fig. 3. Then, based on the above assumption, we bound the difference between the unlabeled loss with pseudo-label $\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ and give the following proposition.
@@ IV-B Generalization Bound for Reward Model
The proof of Theorem 1 can be found in Appendix B. Theorem 1 indicates that the expected error $\mathcal{L}_{R}(\psi)$ is bounded by empirical error $\widehat{\mathcal{L}}_{R}(\psi)$, pseudo-labeling error $\eta\Omega$, Rademacher complexity and constant terms. The constant term can become small when the number of unlabeled data increases, but it may cause unstable training when labels are inaccurate. According to Proposition 1, empirical error and pseudo-labeling error can be reduced through the selection mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$, thus, a better generalization ability, that is tighter upper bound, can be achieved. Moreover, Theorem 1 has generality and can apply to methods that train a reward model using pseudo-labels.
@@ V Policy Learning
To solve the problem of high cost for labeling preference, we propose a offLine prEference-bAsed RL with high Sample Efficiency algorithm (LEASE). This section aims to describe the details of LEASE and establish the theory for a safe policy improvement guarantee. In this paper, we directly use the proposed offline RL algorithms, such as CQL [4] and IQL [25], to perform policy learning, and do not improve the offline RL algorithm itself.
@@ V-A The Implementation Details of LEASE
Algorithm 1 gives pseudo-code for LEASE. The goal of LEASE is to learn a better reward model $R_{\psi}$ from a smaller labeled preference dataset $\mathcal{D}_{l}$ and learn a better policy $\pi_{\theta}$ from an offline dataset $\mathcal{D}_{\text{offline}}$ without interaction with the environment. The key to LEASE is how to train the reward model in the policy learning process. The proposed algorithmic framework comprises three key components: reward model training, transition model training, and policy training. Initially, the transition model is trained using limited offline datasets, while the reward model undergoes pretraining with a small-scale preference dataset. The reward update condition is triggered when the number of unlabeled data in the buffer reaches its maximum capacity. Subsequently, the policy training phase commences. During this phase, the transition model is periodically employed to generate unlabeled synthetic data, and the reward model undergoes iterative updates to enhance its performance.
@@ V-B Safe Policy Improvement Guarantee
Offline RL aims to guarantee $\xi$-safe policy improvement over the behavior policy $\mu$ (the policy used to collect offline dataset), i.e. $J(\widehat{\pi},R^{*})\geq J(\mu,R^{*})-\xi$ [4]. For offline PbRL, we further investigate whether similar policy improvement guarantees can be established. Therefore, this part aims to develop a theoretical guarantee of policy improvement for LEASE, that is, giving the bound for $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$. Then, based on the bounds of policy improvement, we further analyze the sample complexity for LEASE when the reward model is unknown.
@@
The proof of Theorem 2 can be found in Appendix C. The bounds of the PbRL policy learned by LEASE comprise two distinct components: the first term $\xi$ stems from the inherent gap of the offline RL itself, and the second arises from the discrepancy induced by the reward model. The first term $\xi$ is constant when the offline algorithm is determined. The second term is determined by the performance of the reward model. Therefore, the tighter bound can be achieved through reducing the performance gap caused by the reward model (see Proposition 3). It can be reduced to a small value $\varpi$ with a sample complexity of
@@ VI Experiments
  • $\mathbf{Q_{3}}$: How does the proposed selection mechanism and parameter settings affect the performance of reward and policy?
@@ VI-A1 Experimental Datasets
The offline dataset and preference dataset of experiments originate from [22] and [21], respectively. The offline datasets are based on the D4RL dataset [22]. We select medium and medium-expert two types of datasets for Mujoco tasks and human and expert for Adroit tasks. The difference between datasets in a certain task lies in the policies they collect. The Mujoco tasks include halfcheetah-v2, hopper-v2, and walker2d-v2. The Adroit tasks include pen-v1, door-v1, and hammer-v1. Fig. 2 depicts the six tasks for the D4RL benchmark. For three mujoco tasks, the goal is to control the robot’s various joints to achieve faster and more stable locomotion. For adroit tasks, they involve controlling a 24-DoF simulated Shadow Hand robot on a sparse reward, high-dimensional robotic manipulation task.
The preference dataset is from the previous work [21], where they provided the baseline for offline PbRL based on $2000$ preference data labeled by two types (We denote this method as URLHF). One is crowd-sourced labels obtained by crowd-sourcing, and the other is synthetic labels based on ground truth reward. To test LEASE performance for two types of preference data, the labels of preference are from human-realistic feedback for MuJoco tasks, and ground truth reward for Adroit tasks. This paper aims to test the performance under fewer preference data without online interaction. We take the first $100$ data of [21] as $2000$ preference dataset, and denote the method without data augmentation as FEWER.
@@ VI-A2 Experimental Details
Reward and transition model training: Similar to [37], we train an ensemble of $7$ transition models and select the best $5$ models. Each model consists of a $4$-layer feed-forward neural network with $200$ hidden units. The model training employs maximum likelihood estimation with a learning rate of $1\times 10^{-3}$ and Adam optimizer. Following [21], we train an ensemble of $3$ reward models. Each reward model includes a $3$-layer feed-forward neural network with $256$ hidden units. The reward model training uses cross-entropy loss with a learning rate of $3\times 10^{-4}$ and Adam optimizer.
@@
TABLE II: Comparison results of pseudo-labeling accuracy between using the selection mechanism (Ours), only using confidence (No-U), only using uncertainty (No-C), and not using the selection mechanism (No-C-U).
Task Name Ours No-U No-C No-C-U
pen-expert $\mathbf{87.3\pm 0.5}$ $85.9\pm 0.8$ $86.7\pm 0.6$ $86.8\pm 0.5$
door-expert $\mathbf{89.3\pm 0.6}$ $87.8\pm 0.2$ $86.0\pm 0.0$ $85.9\pm 0.8$
Average $\mathbf{87.4\pm 0.4}$ $86.0\pm 0.5$} $85.1\pm 0.6$} $85.4\pm 0.6$}
TABLE III: Results for the D4RL tasks during the last 5 iterations of training averaged over 3 seeds. $\pm$ captures the standard deviation over seeds. Bold indicates the performance within 2% of the best-performing algorithm for each offline algorithm.

### v3

@@ I Introduction
Reinforcement learning (RL) has been widely applied in robot control [1, 10] and healthcare [12, 11] fields. However, in the real-world scenarios, RL faces two serious challenges: 1) Since RL learns the optimal policy through trial and error, it is dangerous for agent to interact directly with environment, especially human-in-loop control [39, 50]; 2) Due to the complexity of human cognitive and control system, it is difficult to design a quantitative reward function to accurately reflect the human intention or preference [9]. Much research has been conducted to solve the above two challenges.
@@ II Related Works
This section provides a brief overview of the algorithms and related theories in offline PbRL, and discusses the differences and advantages between our work and previous associated works.
The algorithm for offline PbRL. Offline PbRL eliminates the demand for interaction with the environment and handcrafted-designed rewards. OPAL [34] is the first algorithm that combines offline RL and PbRL. PT [17] utilized a transformer-based architecture to design a preference model capable of generating non-Markovian rewards. OPPO [30] directly optimized the policy in a high-level embedding space without learning a separate reward function. However, the above methods demand a large amount of preference data. FTB [33] trains a diffusion model to achieve data augmentation, but this consumes a larger training cost. Instead, our method LEASE trains a simple transition model to generate unlabeled data and designs a selection mechanism to ensure the quality of the generated data, which can achieve superior performance under fewer datasets and less time.
The aforementioned approach first learns an explicit reward function from preference data and then utilizes it to guide policy learning. However, in complex tasks, the reward function may contain substantial errors, which can degrade the performance of policy learning. Recently, preference-based direct policy learning has gained significant research attention [51, 52, 53, 54], as it directly optimizes the policy based on user preferences. The most representative work in this area is the Direct Preference Optimization (DPO) algorithm [51], which formulates an initial preference optimization objective. Subsequent research has focused on designing different forms of preference optimization functions to further improve policy learning performance, including simpler preference optimization formulations [52, 53] and optimization objectives based on prospect theory [54]. This direction constitutes a key focus of preference-based policy optimization.
@@ III-A Preliminaries for RL
Preference-based reinforcement learning. Different from the standard RL setting, the reward function is not available in PbRL. Instead, PbRL learns the reward function $\widehat{R}$ from preferences between pairs of trajectory segments to align the human intention [15], where a trajectory segment $\sigma$ of length $L$ is defined as a sequence of states and actions $\{s_{k},a_{k},...,s_{k+L-1},a_{k+L-1}\}\in(\mathcal{S}\times\mathcal{A})^{L}$. Given a pair of segments $(\sigma_{0},\sigma_{1})$, human choose which segment is preferred, i.e., $y\in\{0,1\}$ [14]. The preference label $y=1$ indicates $\sigma_{1}\succ\sigma_{0}$, while $y=0$ indicates $\sigma_{0}\succ\sigma_{1}$, where $\sigma_{i}\succ\sigma_{j}$ denotes that segment $i$ is preferred over segment $j$.
@@
Conversely, given a learned reward model, the preference label between a pair of trajectories ($\sigma_{0},\sigma_{1}$) is determined as follows: if the probability computed by Eq. (2) exceeds 0.5, trajectory 0 ($\sigma_{0}$) is inferred to be preferred over trajectory 1 ($\sigma_{1}$); otherwise, preference is assigned to $\sigma_{1}$. After obtaining the reward function $\widehat{R}$, offline PbRL optimizes the policy by maximizing the expected discounted return $J(\pi,R)$ like standard offline RL.
@@
In practical implementation, to account for the inherent uncertainty in the next-state prediction, a Gaussian distribution is employed to approximate the transition model. To further mitigate the uncertainty associated with model predictions, an ensemble of $N_{T}$ transition models $\{T^{i}_{\varphi}=\mathcal{N}(\mu^{i}_{\varphi},\sigma^{i}_{\varphi})\}_{i=1}^{N_{T}}$ is trained. These models are then optimized using a maximum likelihood function [37]. The samples are generated through $H$-step rollouts. Here, we use the trained transition to generate more trajectories to achieve data augmentation.
@@ III-B Problem Formulation
This assumption states that there must exist a reward function $R\in\mathcal{R}$ which can well-approximate the optimal reward function when the preference data can cover all data space. However, preference data tend to be limited, with even fewer than a hundred in real-world scenarios. Therefore, different from previous PbRL [3, 35, 33], this paper aims to study how to learn a better reward function from the limited dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$. To broaden the coverage of preference data, motivated by model-based offline RL methods, we train the transition model through Eq. (4) to generate more unlabeled preference data $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$. Then, the pseudo labels $\widehat{y}$ for unlabeled dataset $\mathcal{D}_{u}$ can be obtained through reward model trained on $\mathcal{D}_{l}$ [3], that is
$\widehat{y}\left(\sigma_{0}^{u},\sigma_{1}^{u}\right)=\mathbbm{1}\big[P\left(\sigma_{0}^{u}\succ\sigma_{1}^{u};{\psi}\right)<0.5\big],$ (5)
where $\mathbbm{1}(\cdot)$ is indicator function. When the probability is less than 0.5, it indicates that $\sigma_{1}\succ\sigma_{0}$ is highly likely to hold, and the preference label is accordingly set to 1; otherwise, it is set to 0. The reward function can be updated through the collected labeled dataset $\mathcal{D}_{l}$ and the generated unlabeled dataset $\mathcal{D}_{u}$. Then, the reward model can be optimized by minimizing
@@ IV Reward Learning
The reward learning involves two stages: pretraining based on collected labeled data $\mathcal{D}_{l}$ and updating based on $\mathcal{D}_{l}$ and unlabeled data $\mathcal{D}_{u}$ during policy learning, which is essentially semi-supervised learning [26]. However, the generated preference labels may contain errors, which can degrade the performance of the reward model. Therefore, this section focuses on designing a data selection function $f(\sigma_{0}^{u},\sigma_{1}^{u})$ to ensure the quality of the generated preference data and explains the factors influencing the generalization capability of the reward model.
@@ IV-A Uncertainty-aware Pseudo-label Selection
where $\kappa_{p}$ and $\kappa_{\tau}$ are thresholds of confidence and uncertainty respectively. Here, $f(\sigma_{0}^{u},\sigma_{1}^{u})=1$ denotes that the generated data pair $(\sigma_{0}^{u},\sigma_{1}^{u})$ is selected for inclusion, whereas $f(\sigma_{0}^{u},\sigma_{1}^{u})=0$ indicates its exclusion.
Intuitively, more rigorous data selection criteria, characterized by higher confidence and lower variance, lead to a reduction in the number of qualified samples. While such a strategy improves the quality of the generated data, it could lead to a narrower data distribution, which might adversely affect the precision of the trained reward model. Therefore, an appropriate selection of parameters $\kappa_{p}$ and $\kappa_{\tau}$ is required to enhance the performance of the reward model. The unlabeled dataset selected through $f(\sigma_{0}^{u},\sigma_{1}^{u})$ is denoted as $\widetilde{\mathcal{D}}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{\widetilde{N}_{u}}$, where $\widetilde{N}_{u}$ is the number of generated dataset after screening. Then, combining Eq. (6) and the designed selection mechanism, the reward model can be optimized by minimizing
@@
Thanks to its capacity to effectively filter out samples with low confidence and high uncertainty, the screening mechanism mitigates the adverse impact of negative samples on reward model training, thereby contributing to improved reward model performance. Consequently, the bound on the pseudo-label error rate $\eta$ for the filtered generated data is significantly lower than that for the unfiltered data. This observation is experimentally validated in Table III and Fig. 3. Then, based on the above assumption, we bound the difference between the unlabeled loss with pseudo-label $\widehat{\mathcal{L}}_{u}(\psi)$ and that with true label $\widehat{\mathcal{L^{\prime}}}_{u}(\psi)$ and give the following proposition.
@@ IV-B Generalization Bound for Reward Model
The proof of Theorem 1 can be found in Appendix B. Theorem 1 indicates that the expected error $\mathcal{L}_{R}(\psi)$ is bounded by the empirical error $\widehat{\mathcal{L}}_{R}(\psi)$, the pseudo-label error $\eta\Omega$, the Rademacher complexity, and a constant term. While the constant term diminishes as the amount of unlabeled data increases, inaccurate labels can lead to training instability. According to Proposition 1, a selection mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$ can be employed to reduce both the empirical error and the pseudo-label error, thereby achieving better generalization performance, that is, a tighter upper bound. Conversely, without such a screening mechanism, the pseudo-label error rate would rise, resulting in a larger upper bound for the expected error of the reward model and ultimately degrading its performance. Moreover, Theorem 1 has generality and can apply to methods that train a reward model using pseudo-labels.
@@ V Policy Learning
To solve the problem of high cost for labeling preference, we propose a offLine prEference-bAsed RL with high Sample Efficiency algorithm (LEASE). This section aims to describe the details of LEASE and establish the theory for a safe policy improvement guarantee. In this paper, we directly use the proposed offline RL algorithms, such as Conservative Q-Learning (CQL) [4] and Implicit Q-Learning (IQL) [25], to perform policy learning, and do not improve the offline RL algorithm itself.
@@ V-A The Implementation Details of LEASE
Algorithm 1 gives pseudo-code for LEASE. The goal of LEASE is to learn a better reward model $R_{\psi}$ from a smaller labeled preference dataset $\mathcal{D}_{l}$ and learn a better policy $\pi_{\theta}$ from an offline dataset $\mathcal{D}_{\text{offline}}$ without interaction with the environment. The key to LEASE is how to train the reward model in the policy learning process. The proposed algorithmic framework comprises three key components: reward model training, transition model training, and policy training. Initially, the transition model is trained on a limited offline dataset, whereas the reward model is pretrained using a small-scale preference dataset. The reward update mechanism is activated once the volume of unlabeled data in the buffer reaches its upper capacity limit. Subsequently, the policy training phase is initiated. In this phase, the transition model is periodically leveraged to generate unlabeled synthetic data, and the reward model undergoes iterative updates to further boost its performance.
@@ V-B Safe Policy Improvement Guarantee
Offline RL aims to guarantee $\xi$-safe policy improvement over the behavior policy $\mu$ (the policy used to collect offline dataset), i.e. $J(\widehat{\pi},R^{*})\geq J(\mu,R^{*})-\xi$ [4]. For offline PbRL, we further investigate whether similar policy improvement guarantees can be established. Unlike offline RL, the reward function in PbRL is unknown, and the learned reward function may contain errors, which introduces an additional performance gap. Therefore, this part aims to develop a theoretical guarantee of policy improvement for LEASE, that is, giving the bound for $J(\mu,R^{*})-J(\widehat{\pi},R^{*})$. Then, based on the bounds of policy improvement, we further analyze the sample complexity for LEASE when the reward model is unknown.
@@
The proof of Theorem 2 can be found in Appendix C. The performance bound for the PbRL policy learned by LEASE consists of two distinct components: the first term, denoted as $\xi$, stems from the inherent gap of the offline RL process itself; the second term arises from the discrepancy induced by the reward model. When the offline algorithm is fixed, the first term $\xi$ is constant. The second term depends on the performance of the learned reward model. Consequently, the proposed framework is readily compatible with other offline RL algorithms, and a tighter overall bound can be achieved by reducing the performance gap attributable to the reward model (see Proposition 3). It can be reduced to a small value $\varpi$ with a sample complexity of
@@ VI Experiments
  • $\mathbf{Q_{3}}$: How do the proposed selection mechanism and parameter settings affect the performance of reward and policy?
@@ VI-A1 Experimental Datasets
The offline dataset and preference dataset of experiments originate from [22] and [21], respectively. The offline datasets are based on the D4RL dataset [22]. We select two types of datasets for Mujoco tasks: medium and medium-expert, and human and expert for Adroit tasks. The difference between datasets in a certain task lies in the policies they collect. The Mujoco tasks include halfcheetah-v2, hopper-v2, and walker2d-v2. The Adroit tasks include pen-v1, door-v1, and hammer-v1. Fig. 2 depicts the six tasks for the D4RL benchmark. The objective of the three Mujoco tasks is to control the robot’s joints to achieve faster and more stable motion, with action dimensions ranging from 3 to 8, and the reward function is defined in the form of continuous values. In contrast, the Adroit task involves controlling a simulated Shadow Hand robot with an action dimension of 24, which is a high-dimensional robotic manipulation task characterized by sparse rewards—only a few specific reward values are provided.
The preference dataset is from the previous work [21], where they provided the baseline for offline PbRL based on $2000$ preference data labeled by two types (We denote this method as URLHF). One is crowd-sourced labels obtained through crowd-sourcing, and the other is synthetic labels based on ground-truth rewards. To test LEASE performance for two types of preference data, the labels of preference are from human-realistic feedback for MuJoco tasks, and ground truth reward for Adroit tasks. This paper aims to test the performance under fewer preference data without online interaction. We take the first $100$ data of [21] as $2000$ preference dataset, and denote the method without data augmentation as FEWER.
@@ VI-A2 Experimental Details
Reward and transition model training: Similar to [37], we train an ensemble of $7$ transition models and select the best $5$ models. Each model consists of a $4$-layer feed-forward neural network with $200$ hidden units. The model training employs maximum likelihood estimation with a learning rate of $1\times 10^{-3}$ and Adam optimizer. Following [21], we train an ensemble of $3$ reward models. Each reward model includes a $3$-layer feed-forward neural network with $256$ hidden units. The training of the reward model uses cross-entropy loss with a learning rate of $3\times 10^{-4}$ and the Adam optimizer.
@@
TABLE II: Results for the D4RL tasks during the last 5 iterations of training averaged over 3 seeds. $\pm$ captures the standard deviation over seeds. Bold indicates the performance within 2% of the best-performing algorithm for each offline algorithm.